# Domain decomposition of the modified Born series approach for large-scale wave propagation simulations

### v1

@@ 4.3 Large scale simulations
To demonstrate the larger simulation size enabled by the domain decomposition approach, we run a simulation of $3.1\cdot 10^{7}$ cubic wavelengths on a system with an Intel Xeon Silver 4216 CPU @ 2.10 GHz, 256 GB RAM, and two GPUs (both NVIDIA A40 48GB). The structure is the same as mentioned at the beginning of Section 4, but the size is a much larger $315\!\times\!315\!\times\!$ 315 × 315 wavelengths, corresponding to $1260\!\times\!1260\!\times\!$ 1260 × 1260 voxels. Due to this large size, the simulation cannot be performed on a single GPU. Therefore, we run the single-domain simulation on the CPU and compare it against the 2-domain, 2-GPU simulation performed with the same system.
Fig. 8a shows a slice of the 3D random refractive index map that was used. The results of the single-domain and 2-domain simulation are shown in Figs. 8b and c, respectively. The relative error between the two is $5.2\cdot 10^{-6}$.
Figure 8: Large 3D simulation of size $315\!\times\!315\!\times\!$ 315 × 315 wavelengths with and without domain decomposition. A 2D slice of (a) the random refractive index, (b) the result from a 1-domain simulation on the CPU, and (c) the result from a 2-domain simulation using two GPUs, with the dashed grey lines indicating the subdomains. (d) Residual as a function of the iterations for (b), a solid line, and (c), a dashed line.
Fig. 8d shows the residual as a function of the iterations for the two simulations. It can be seen that the residual decreases monotonically, which is a key property of the MBS. The two-domain simulation converges in 657 iterations in 6.3 minutes (379 s), while the single-domain simulation converges in 343 iterations in 12 hours (43179 s) since it is run on the CPU (Table 1).
In the 2-domain case, the wrapping and transfer corrections give rise to a lower scaling coefficient $c$, causing the number of iterations to increase by 1.9×1.9$\times$1.9 ×. In addition to this increase in the number of iterations, two overheads are associated with the domain decomposition framework. First is the overhead of computing the edge corrections. Second, the overhead due to data transfer between GPUs when the edge corrections are transferred and applied to neighbouring subdomains in each iteration. This leads to a lock-step execution, meaning the GPUs spend some time waiting for the results of the other GPU(s) to arrive. This overhead adds approximately $40\%$ time per cubic wavelength and reduces for larger simulations.
@@
2 domains, 2 GPUs $3.1\cdot 10^{7}$ $378.63$ $657$
1 domain, CPU $3.1\cdot 10^{7}$ $43179.14$ $343$
Table 1: The largest possible simulation with domain decomposition, using 2 GPUs, and without, using the CPU. The simulations were run on a system with an Intel Xeon Silver 4216 CPU @ 2.10 GHz with 256 GB RAM and two GPUs (both NVIDIA A40 48GB). The domain decomposition simulation (row 1) takes 1.9×1.9$\times$1.9 × the number of iterations for convergence as the CPU simulation (row 2) but is 114×114$\times$114 × faster.
Despite this overhead, the 2-domain, 2-GPU simulation is 114×114$\times$114 × faster than the single-domain simulation on the CPU without domain decomposition. With these two GPUs, we increased the simulation size by 1.93×1.93$\times$1.93 ×, up to an unprecedented $315\!\times\!315\!\times\!$ 315 × 315 wavelengths. Furthermore, we showed that the number of domains can be increased arbitrarily. The overhead of adding additional domains is much smaller than the initial step from one to two subdomains, especially when the domain is split into subdomains along a single axis.
@@ 5 Conclusion
We have introduced a domain decomposition of the modified Born series (MBS) approach [2, 34] applied to the Helmholtz equation. With the new framework, we simulated a complex 3D structure of a remarkable $3.1\cdot 10^{7}$ wavelengths in size in just $379$ seconds by solving over two GPUs. This represents a factor of $1.93$ increase over the largest possible simulation on a single GPU without domain decomposition.
Our decomposition framework hinges on the ability to split the linear system as $A=L+V$. Instead of the traditional splitting, where $V$ is a scattering potential that acts locally on each voxel, we introduced a $V$ that includes the communication between subdomains and corrections for wraparound artefacts. As a result, the operator $(L+I)^{-1}$ in the MBS iteration can be evaluated locally on each subdomain using a fast convolution. Therefore, this operator, which is the most computationally intensive step of the iteration, can be evaluated in parallel on multiple GPUs.
Despite the significant overhead of our domain splitting method due to an increased number of iterations, and communication and synchronisation overhead, the ability to split a simulation over multiple GPUs results in a very significant speedup. Already, with the current dual-GPU system, we were able to solve a problem of $315\times 315\times 315$ wavelengths 114×114$\times$114 × faster than without domain decomposition since the non-decomposed problem is too large to fit on a single GPU. Moreover, as we demonstrated in Section 4.2.2, there is only a little overhead associated with adding more subdomains along an axis after the first splitting. This favourable scaling paves the way for distributing simulations over more GPUs or compute nodes in a cluster.

### v2

@@ 4.3 Large scale simulations
To demonstrate the larger simulation size enabled by the domain decomposition approach, we run a simulation of $3.1\cdot 10^{7}$ cubic wavelengths on a system with an AMD EPYC 9354P CPU @ 3.25 GHz with 256 GB RAM and two GPUs (both NVIDIA L40S 48GB). The structure is the same as mentioned at the beginning of Section 4, except that the spheres have a radius of $6$ wavelengths and the clearance between them is set to $6$ wavelengths, which helps the energy propagate deeper into the simulation away from the source. The simulation size is a much larger $315\times 315\times 315$ wavelengths, corresponding to $1260\times 1260\times 1260$ voxels. Due to this large size, the simulation cannot be performed on a single GPU. Therefore, we run the single-domain simulation on the CPU and compare it against the 2-domain, 2-GPU simulation performed with the same system.
Fig. 8a shows a slice, through the centre of the z-axis, of the 3D refractive index map used. The single-domain and 2-domain simulation results are shown in Figs. 8b and c, respectively. The insets in Fig. 8a-c show finer details in a zoomed-in region of the slices. The relative error between the two is $2.5\cdot 10^{-4}$.
Figure 8: Large 3D simulation of size $315\times 315\times 315$ wavelengths with and without domain decomposition. A 2D slice of (a) the refractive index map, (b) the result from a 1-domain simulation on the CPU, and (c) the result from a 2-domain simulation using two GPUs, with the dash-dotted black line indicating the subdomain boundary. The insets are zoomed-in portions of the selected area to show finer details. (d) Residual as a function of the iterations for (b), a solid line, and (c), a dashed line.
Fig. 8d shows the residual as a function of the iterations for the two simulations. The two-domain simulation converges in 9,878 iterations in 1.4 hours (4,958 s), while the single-domain simulation converges in 1,181 iterations in 18.2 hours (65,537 s) since it is run on the CPU (Table 1).
In the 2-domain case, the wrapping and transfer corrections lower the scaling coefficient $c$, increasing the number of iterations by 8.3×8.3$\times$8.3 ×. In addition to this increase in the number of iterations, the domain decomposition framework has two overheads: the overhead of computing the edge corrections and the overhead due to data transfer between GPUs when the edge corrections are transferred and applied to neighbouring subdomains in each iteration. This subdomain communication leads to a lock-step execution, meaning the GPUs wait for the results of the other GPU(s). This overhead adds approximately $40\%$ time per cubic wavelength and reduces for larger simulations.
@@
2 domains, 2 GPUs $3.1\cdot 10^{7}$ $4,958.65$ $9,878$
1 domain, CPU $3.1\cdot 10^{7}$ $65,537.16$ $1,181$
Table 1: The largest possible simulation with domain decomposition, using 2 GPUs, and without, using the CPU. The simulations were run on a system with an AMD EPYC 9354P CPU @ 3.25 GHz with 256 GB RAM and two GPUs (both NVIDIA L40S 48GB). The domain decomposition simulation (row 1) takes 8.3×8.3$\times$8.3 × the number of iterations for convergence as the CPU simulation (row 2) but is 13.2×13.2$\times$13.2 × faster.
Despite this overhead, the 2-domain, 2-GPU simulation is 13.2×13.2$\times$13.2 × faster than the single-domain simulation on the CPU without domain decomposition. With these two GPUs, we increased the simulation size by 1.93×1.93$\times$1.93 × of what can be accommodated on a single GPU to an unprecedented $315\times 315\times 315$ wavelengths. Furthermore, we showed that the number of domains can be increased arbitrarily. The overhead of adding additional domains is much smaller than the initial step from one to two subdomains, especially when the domain is split into subdomains along a single axis.
@@ 5 Conclusion
We have introduced a domain decomposition of the modified Born series (MBS) approach [2, 40] applied to the Helmholtz equation. With the new framework, we simulated a complex 3D structure of a remarkable $3.1\cdot 10^{7}$ wavelengths in size in just $1.4$ hours by solving over two GPUs. This result is a factor of $1.93$ increase over the largest possible simulation on a single GPU without domain decomposition.
Our decomposition framework hinges on the ability to split the linear system as $A=L+V$. Instead of the traditional splitting, where $V$ is a scattering potential that acts locally on each voxel, we introduced a $V$ that includes the communication between subdomains and corrections for wraparound artefacts. As a result, the operator $(L+I)^{-1}$ in the MBS iteration can be evaluated locally on each subdomain using a fast convolution. Therefore, this operator, the most computationally intensive step of the iteration, can be evaluated in parallel on multiple GPUs.
Despite the significant overhead of our domain splitting method due to an increased number of iterations and communication and synchronisation overhead, the ability to split a simulation over multiple GPUs results in a significant speed-up. Already, with the current dual-GPU system, we were able to solve a problem of $315\times 315\times 315$ wavelengths 13.2×13.2$\times$13.2 × faster than without domain decomposition since the non-decomposed problem is too large to fit on a single GPU. Moreover, as we demonstrated in Section 4.2.2, there is only a slight overhead associated with adding more subdomains along an axis after the first splitting. This favourable scaling paves the way for distributing simulations over more GPUs or compute nodes in a cluster.